A 0.825 mol sample of NO2(g) initially at 298 K and 1.00 atm is held at constant volume while enough heat is applied to raise the temperature of the gas by 19.3 K.

Assuming ideal gas behavior, calculate the amount of heat (?) in joules required to affect this temperature change and the total change in internal energy, Δ?. Note that some books use Δ? as the symbol for internal energy instead of Δ?.
Answer #1

Since the process is carried out at constant volume (\DeltaV=0), work done, w = - P\DeltaV =0

The heat supplied is completely utilized to increase the internal energy of the system and is given by

q = \DeltaU = nCv\DeltaT =n*3R\DeltaT = 0.825*3*8.314*(19.3-298)= -5734.85 joules.

i.e 5734.85 joules of heat is taken from the system.